# Efficient Multilingual Search Relevance Modeling in E-Commerce via LLM Mixture-of-Experts

### v1

@@ Abstract.
In e-commerce platforms, search relevance directly influences both user experience and merchant revenue. In multi-country deployments, diverse linguistic, cultural, and product catalog contexts introduce significant distribution shifts, posing substantial challenges to relevance modeling. Existing approaches typically enhance the reasoning or multilingual abilities of a single monolithic model, yet they remain limited by data diversity, coverage gaps, and high inference costs in heterogeneous environments. Our empirical analysis reveals that different LLM base models exhibit complementary strengths across languages and regions, motivating an expert-based architecture. We propose a scalable LLM-based Mixture-of-Experts (MoE) framework that dynamically routes queries to specialized experts and fuses their embeddings through concatenation. Among rule-based, pseudo-label–based, and fully end-to-end strategies, end-to-end hard routing with concatenation offers the best balance of effectiveness and efficiency. To mitigate inference overhead, we further develop an engineering-optimized offline batch pipeline with resource-efficient scheduling, which hides memory latency, improves GPU utilization, and reduces GPU-hour consumption by up to 35% compared with synchronous execution. On datasets spanning six Southeast Asian markets, our MoE improves AUC by 0.72 percentage points over a dense baseline with the same active parameters. Meanwhile, the optimized pipeline achieves 27.6 queries per second (QPS), a 9% throughput improvement. These results demonstrate superior multilingual relevance and efficiency, delivering strong cost-effectiveness for real-world e-commerce search systems.
@@
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: The ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesjournalyear: 2026isbn: 978-1-4503-XXXX-X/2018/06submissionid: 210ccs: Applied computing Online shoppingccs: Computing methodologies Natural language generationccs: Information systems Query intent
Figure 1. Radar chart showing AUC of three base models across multiple country–language tasks. :
@@ 1. Introduction
In recent years, the rapid growth of e-commerce platforms across multiple countries has made search relevance in multilingual and multicultural environments a critical factor affecting both user experience and merchant revenue. Traditional approaches typically employ BERT-like models to learn contextualized representations of queries and product descriptions, determining relevance via semantic similarity. With the advent of large language models (LLMs), their extensive knowledge and strong language understanding have opened a new paradigm for relevance estimation.
@@
Most existing LLM-based methods focus on enhancing the capabilities of a single model, such as improving reasoning through fine-tuning (10.1145/3701716.3715246; zhao2025explainable), incorporating multimodal signals (Zheng_2023; zhu2023querylifequeryawarelanguageimage) like product images or user behavior, or performing continued pre-training (dou2025sailor2sailingsoutheastasia) to strengthen multilingual understanding. While these approaches can achieve notable gains, they often demand substantial training and inference resources, and collecting sufficiently diverse and representative data across multiple real-world markets remains challenging. Such single-model enhancements can be directly integrated into our proposed framework as independent experts. Our empirical analysis (Figure 1) shows that even with identical fine-tuning data/objectives, different LLMs (Gemma2-9B (gemma_2024), Qwen2.5-14B (qwen2.5), Sailor2-20B (dou2025sailor2sailingsoutheastasia)) vary in performance across languages/regions. For example, Sailor2-20B (larger scale, Southeast Asian languages pre-training) does not outperform others in some languages, while Qwen2.5-14B excels in Indonesian–English and Singapore–Vietnamese. This suggests that differences in pre-training corpora, vocabulary design, and language coverage lead each model to internalize distinct regional language capabilities, which remain complementary even under identical fine-tuning.
@@
Building on these observations, we propose a lightweight LLM-based Mixture-of-Experts (MoE) framework that integrates fine-tuned LLMs as independent experts without additional fine-tuning. The routing module dynamically assigns each input to the most suitable experts, and their representations are fused at the embedding level, leveraging both the improvements within each expert and the complementary strengths across models. This design exploits the advantages of multiple base models in multilingual and multi-regional relevance tasks, achieving a balanced trade-off between effectiveness and efficiency. We explore various routing strategies, including rule-based, pseudo-label, and end-to-end (soft/hard) approaches, and find that end-to-end hard routing delivers the best overall balance. Experiments on real-world multi-country e-commerce datasets show that our method improves AUC by 0.6 percentage points and throughput by approximately 120% over a strong dense baseline, while maintaining low inference latency.
@@ 2. Methodology
To leverage the complementary strengths of LLMs in multilingual e-commerce search while controlling inference cost, we propose a sparsely-activated Mixture-of-Experts (MoE) framework—where each LLM acts as an independent expert. This framework jointly optimizes expert selection and representation fusion to balance high relevance performance with predictable computational overhead.
@@
2.1. LLM-based Mixture-of-Experts Architecture
Our framework integrates multiple finetuned LLMs as experts. Given a query–item pair, a routing module first selects a subset of experts, whose contextual embeddings are then fused for final relevance prediction. The routing module balances prediction performance and computational efficiency. We consider four strategies:
@@ 2.1. LLM-based Mixture-of-Experts Architecture
End-to-end soft routing outputs a probability distribution over all $N$ experts and applies an entropy minimization regularizer to encourage sparsity: $L_{\text{ent}}=\sum_{i=1}^{N}p_{i}\log p_{i}$, where $p_{i}$ is the routing probability for expert $i$. This encourages sharper expert selection and model specialization. During inference, only experts exceeding a threshold are activated. Soft routing may suffer from threshold sensitivity and train–inference mismatch.
End-to-end hard routing directly selects the top-$k$ experts (e.g., $k=2$) and incorporates a load balancing loss to prevent routing collapse. Differences in pre-training corpora and architectures produce heterogeneous hidden representations across experts. Early in training, the fusion module may overfit to the easiest expert, causing vanishing gradients for the router. To mitigate this, we adopt a load balancing loss (fedus2022switchtransformersscalingtrillion):
@@
Table 1. Overall and per-market AUC (%) and QPS for all baselines and our proposed MoE. Bold indicates the best, underline the second-best.
@@
For each selected expert $i$, we extract its final transformer layer’s last-token hidden state $\mathbf{h}_{i}\in\mathbb{R}^{d_{i}}$, then project it to a shared space:
(2) $\mathbf{h}^{\prime}_{i}=W_{i}\mathbf{h}_{i}+\mathbf{b}_{i},\quad\mathbf{h}^{\prime}_{i}\in\mathbb{R}^{d}.$
The projected embeddings are then concatenated:
(3) $\mathbf{h}_{\text{concat}}=[\mathbf{h}^{\prime}_{i_{1}}\parallel\dots\parallel\mathbf{h}^{\prime}_{i_{k}}]\in\mathbb{R}^{kd}.$
Concatenation preserves non-redundant knowledge (e.g., syntax, cultural context) from multiple experts, unlike weighted fusion which may prematurely compress features. The fused vector is passed through a lightweight classifier:
(4) $\hat{y}=\sigma(W_{c}\mathbf{h}_{\text{concat}}+\mathbf{b}_{c}).$
The training loss combines cross-entropy and load balancing:
(5) $L=L_{\text{CE}}+\lambda L_{\text{LB}},$
where $\lambda$ balances task performance and expert diversity.
@@ 3.2. Experimental Settings
Base models—Qwen2.5-14B, Gemma2-9B, Sailor2-20B, and Qwen2.5-32B—are fine-tuned on the same training set using RSLora(kalajdzievski2023rankstabilizationscalingfactor) with rank $r=256$ and scaling factor $\alpha=128$. Training is conducted for a single epoch with batch size 4 and learning rate $5\times 10^{-6}$, following standard practice for large-scale relevance fine-tuning. The MoE framework, including the router, projection layers, and fusion classifier, is trained with batch size 128, learning rate $1\times 10^{-4}$, and $\lambda=0.01$ in Eq. 5, while all base LLMs remain frozen.
@@ 3.5. Online Deployment
In industrial search advertising systems, upon receiving a user query, a Top-K subset of ads is selected from a vast inventory, following a standard multi-stage pipeline: index → retrieval → prerank → rank → auction → rerank. At the retrieval stage, query–item relevance scores serve dual purposes: filtering candidates for downstream stages and informing dynamic reserve prices in the auction mechanism (feng2018deep; Li_2025), which balance user experience, platform revenue, and advertiser ROI. Accurate relevance estimation is therefore critical for both result quality and monetization efficacy.
However, industrial-scale retrieval involves tens to hundreds of thousands of candidate items per query. Real-time inference with large language models (LLMs) at this scale is computationally prohibitive due to latency constraints. To leverage LLM semantics without losing efficiency, we use offline knowledge distillation: LLMs generate high-quality relevance labels on a billion-scale query-item corpus to train a compact student model. We select ColBERT (khattab2020colbertefficienteffectivepassage) as the student model due to its strong representational capacity and efficiency in late-interaction architectures. The distilled ColBERT model is deployed online, enabling fine-grained, low-latency relevance scoring that meets the demands of real-time retrieval.
It should be noted that the translation of improved AUC into actual gains in ad revenue is influenced by multiple confounding factors, such as dynamic reserve price parameters and shifts in the online bidding environment. To ensure rigorous interpretation and avoid potential misattribution, we refrain from reporting online experimental results at this stage.

### v2

# Orchestrating Heterogeneous Experts: A Scalable MoE Framework with Anisotropy-Preserving Fusion

@@ Abstract.
In cross-border e-commerce, search relevance modeling faces the dual challenge of extreme linguistic diversity and fine-grained semantic nuances. Existing approaches typically rely on scaling up a single monolithic Large Language Model (LLM). However, our empirical analysis reveals that single models suffer from uneven capability distributions across regions. For example, excelling in English while underperforming in specific Southeast Asian languages. In this work, we shift the paradigm from scaling a single model to orchestrating heterogeneous experts. We propose a scalable Coarse-grained Mixture-of-Experts (MoE) framework that leverages the inherent complementarity of distinct open-source LLMs (e.g., Qwen, Gemma) without expensive pre-training. Unlike standard token-level MoE, our framework dynamically routes entire queries to specialized experts and, crucially, employs an Information-Preserving Concatenation Fusion strategy. We theoretically posit that preserving the distinct embedding manifolds of heterogeneous experts—rather than compressing them via weighted averaging—is essential for capturing complex relevance signals in a multi-model latent space. On datasets spanning six Southeast Asian markets, our MoE improves AUC by 0.72 percentage points over a dense baseline with the same active parameters. Meanwhile, the optimized pipeline achieves 13.72 queries per second (QPS), a 9% throughput improvement.
@@
copyright: acmcopyrightjournalyear: 2026doi: xxxxx/xxxxxxxconference: Companion Proceedings of the ACM Web Conference 2026,; April 13–17, 2026; Dubai, United Arab Emiratesisbn: 979-8-4007-2308-7/2026/04ccs: Applied computing Online shoppingccs: Computing methodologies Natural language generationccs: Information systems Query intent
Figure 1. Radar chart showing the AUC performance of three base models across multiple country–language tasks. Different models exhibit distinct capability skews (e.g., Qwen excels in ID-en, Sailor in TH-th), highlighting the potential for complementarity through a Mixture-of-Experts approach.
Radar chart showing AUC of three base models across multiple country–language tasks, with curves overlapping and each model leading in different areas.
@@ 1. Introduction
In the rapidly expanding landscape of cross-border e-commerce, search relevance modeling serves as the linchpin for both user experience and platform revenue. Unlike standard web search, e-commerce queries in Southeast Asian markets are characterized by extreme brevity, intent ambiguity, and, crucially, a highly fragmented linguistic distribution. A single platform must often process English queries requiring deep semantic reasoning while simultaneously handling local languages (e.g., Thai, Indonesian) where syntactic exactness and local entity recognition are paramount.
Traditionally, the industry has relied on scaling up monolithic Pre-trained Language Models (PLMs), such as BERT-based encoders or larger unified LLMs, to address this diversity. However, this ”one-model-fits-all” paradigm faces a steep curve of diminishing returns. Training a single massive model to master all linguistic nuances and domain-specific knowledge (e.g., local brands, slang) is not only computationally prohibitive but often leads to the ”curse of multilingualism,” where optimizing for high-resource languages can inadvertently degrade performance on low-resource ones due to capacity dilution (Conneau et al., 2020).
@@
Our empirical analysis (Figure 1) shows that even with identical fine-tuning data/objectives, different LLMs (Gemma2-9B(Team, 2024a), Qwen2.5-14B(Team, 2024b), Sailor2-20B(Dou et al., 2025)) vary in performance across languages/regions. For example, Sailor2-20B (larger scale, Southeast Asian languages pre-training) does not outperform others in some languages, while Qwen2.5-14B excels in Indonesian–English and Singapore–Vietnamese. This suggests that differences in pre-training corpora, vocabulary design, and language coverage lead each model to internalize distinct regional language capabilities, which remain complementary even under identical fine-tuning. This observation drives our core hypothesis: Instead of forcing a single model to learn everything, can we orchestrate a mixture of specialized, heterogeneous experts to achieve a Pareto-optimal balance between performance and cost?
@@
To leverage this complementarity, we propose a lightweight LLM-based Mixture-of-Experts (MoE) framework that integrates fine-tuned LLMs as independent experts without additional fine-tuning. The routing module dynamically assigns each input to the most suitable experts, and their representations are fused at the embedding level, leveraging both the improvements within each expert and the complementary strengths across models. We explore various routing strategies, including rule-based, pseudo-label, and end-to-end (soft/hard) approaches, and find that End-to-end Hard Routing delivers the best overall balance.
Another key design decision in our framework is the embedded fusion strategy, which has often been overlooked in previous work. Traditional ensemble methods typically employ Weighted average (Scalar Mixing) to combine expert outputs. However, we argue that simply averaging embeddings from heterogeneous architectures (with distinct latent manifolds and tokenizer spaces) causes destructive interference. Since the basis vectors of Expert A and Expert B are not naturally aligned, a linear combination often collapses distinct semantic signals into noise. To address this, we introduce an Information-Preserving Concatenation mechanism. By preserving the full dimensionality of expert outputs in a concatenated space, we allow the downstream classifier to learn a non-linear decision boundary, effectively acting as a discriminator that learns which expert to trust for specific feature subspaces (e.g., relying on Qwen for syntax and Gemma for attributes).
Furthermore, to ensure industrial viability, we implement an asynchronous batch inference pipeline, decoupling heavy expert computation from real-time routing logic. This allows us to deploy powerful LLM experts within strict latency constraints.
Our main contributions are summarized as follows.
  • Architectural Insight: We identify the distinct complementarity between heterogeneous LLMs in cross-border e-commerce and propose a framework to exploit this diversity without expensive continued pre-training.
  • Manifold-Preserving Fusion: We theoretically and empirically demonstrate that concatenation-based fusion outperforms traditional scalar mixing by preserving the distinct geometric structures of heterogeneous expert embeddings, resolving the issue of feature misalignment.
  • Industrial Effectiveness: Validated on a large-scale real-world dataset spanning six markets, our approach achieves a 0.72 percentage point AUC gain over parameter-equivalent dense baselines and significantly improves throughput via an optimized asynchronous pipeline.
@@ 2. Methodology
2.1. Framework Overview
To address the challenges of linguistic diversity and semantic complexity in cross-border e-commerce, we propose a scalable Coarse-grained Mixture-of-Experts (MoE) framework. Unlike standard token-level MoE architectures that activate experts for each token generation step, our framework operates at the request level, dynamically routing entire queries to specialized Large Language Models (LLMs) and fusing their semantic representations.
Notation. We use lowercase bold letters (e.g., $\mathbf{x}$$\mathbf{h}$) to denote vectors and uppercase bold letters (e.g., $\mathbf{W}$) for matrices. Sets are denoted by calligraphic letters (e.g., $\mathcal{S}$). $\mathit{N}$ denotes the total number of experts and k denotes the number of active experts per query.
Architecture Overview. As illustrated in Figure  2, the inference pipeline consists of three decoupled stages designed to balance effectiveness and efficiency:
(1) Dynamic Routing: A lightweight routing module analyzes the input tuple (q,t,c) and selects a sparse subset of top-k experts, $\mathcal{S}_{k}\subset\{1,\ldots,N\}$, where $k\ll N$, from a set of N heterogeneous LLM experts. This sparsity ensures that computational resources are allocated only to the most suitable models for a given region or language.
(2) Expert Inference: The selected experts process the input in parallel. For each selected expert $i\in\mathcal{S}_{k}$, we extract the last-token hidden state $\mathbf{h}_{i}$ from its final transformer layer. This vector encapsulates the expert’s semantic understanding of the query-item pair.
[…]
The overall training objective minimizes the discrepancy between $\hat{y}$ and ground truth y, regularized by a load-balancing term to prevent routing collapse. The details of the routing and fusion mechanisms are described in Section 2.2 and Section 2.3, respectively
2.2. Dynamic Sparse Routing
The routing module balances predictive performance and computational efficiency. We considered four strategies:
@@ 2.2. Dynamic Sparse Routing
End-to-end soft routing outputs a probability distribution over all $N$ experts and applies an entropy minimization regularizer to encourage sparsity. This encourages sharper expert selection and model specialization. During inference, only experts exceeding a threshold are activated. Soft routing may suffer from threshold sensitivity and train–inference mismatch.
End-to-end hard routing directly selects the top-$k$ experts (e.g., $k=2$) and incorporates a load balancing loss to prevent routing collapse. Differences in pre-training corpora and architectures produce heterogeneous hidden representations across experts. Early in training, the fusion module may overfit to the easiest expert, causing vanishing gradients for the router. To mitigate this, we adopt a load balancing loss (Fedus et al., 2022):
@@
Table 1. Overall and per-market AUC (%) and QPS for all baselines and our proposed MoE. Bold indicates the best, underline the second-best. Our Proposed MoE achieves the best relevance performance across most markets while maintaining competitive throughput compared to dense baselines. Note that ”Full Fusion” lags behind MoE, validating the effectiveness of sparse routing.
@@
2.3. Manifold-Preserving Fusion Strategy
A crucial challenge in utilizing heterogeneous experts lies in the aggregation of their output. While standard MoE architectures typically employ Scalar Mixing (Weighted Averaging), defined as $\mathbf{h}_{mix}=\sum w_{k}\mathbf{h_{k}}$, we demonstrate that this approach is theoretically suboptimal for aggregating heterogeneous LLMs.
The Misalignment Problem and Destructive Interference. Scalar mixing relies on the strong assumption that the latent spaces of all experts are isomorphic and semantically aligned—i.e., the i-th dimension of Expert A’s embedding vector encodes the same semantic feature as the i-th dimension of Expert B.
However, for heterogeneous models (e.g., Qwen vs. Gemma) pre-trained on distinct corpora with different tokenizers and architectures, the latent manifolds are topologically distinct. Even if projected to the same dimensionality, the basis vectors are not naturally aligned. Consequently, performing a linear combination on these unaligned vectors leads to destructive interference, where a distinct feature signal in one expert (e.g., a specific syntax pattern) may be cancelled out or diluted by noise from another expert in the same dimension. This results in a ”feature collapse” where fine-grained semantic information is lost during aggregation.
Manifold Preservation via Concatenation. To address this, we propose a Manifold-Preserving Fusion strategy. Instead of collapsing the representations into a shared lower-dimensional space, we construct a joint representation space by concatenating the selected expert embeddings.
First, for each selected expert i, we extract its final transformer layer’s last-token hidden state $\mathbf{h}_{i}\in\mathbb{R}^{d_{i}}$ and project it to a normalized space:
(2) $\mathbf{h}_{i}^{\prime}=\mathbf{W}_{i}\mathbf{h_{i}}+\mathbf{b}_{i},\mathbf{h}_{i}^{\prime}\in\mathbb{R}^{d}$
Where $\mathbf{W}_{i}\in\mathbb{R}^{d\times d_{i}}$ projects the expert-specific dimension $\textit{d}_{i}$ to a shared dimension d.
Then, we form the joint representation z by concatenating the projected embeddings:
(3) $\mathbf{z}=[\mathbf{h}_{1}^{\prime};\mathbf{h}_{2}^{\prime};\ldots;\mathbf{h}_{k}^{\prime}]\in\mathbb{R}^{k\times d}$
Geometrically, this operation preserves the intrinsic manifold structure of each expert within independent subspaces of z, avoiding the destructive interference inherent in averaging.
Non-linear Decision Boundary. To extract relevant signals from this preserved high-dimensional space, we employ a light-weight classifier (MLP) as a non-linear discriminator to predict the relevance score $\hat{y}$:
(4) $\hat{y}=\sigma(\mathbf{W_{c}}\cdot\text{ReLU}(\mathbf{W}_{p}\mathbf{z}+\mathbf{b}_{p})+\mathbf{b}_{c})$
Unlike a simple dot-product attention mechanism used in scalar mixing, this learnable MLP allows the model to capture cross-expert interactions (e.g., ”if Expert A detects high semantic relevance AND Expert B detects brand mismatch, then label as negative”). This enables the system to effectively select the most reliable signal source for the specific query instance without suffering information loss.
Training Loss. The model is trained end-to-end. We employ the standard Cross-Entropy loss for the relevance classification task:
[…]
where $\mathcal{B}$ is the training batch. The final objective combines this with the load balancing loss $L_{LB}$ (defined in Eq. 1) to ensure expert diversity:
(6) $\mathcal{L}{total}=L_{CE}+\lambda L_{LB}$
where $\lambda$ is a hyperparameter that balances task performance and expert utilization.
@@ 3.2. Experimental Settings
Base models—Qwen2.5-14B, Gemma2-9B, Sailor2-20B, and Qwen2.5-32B—are fine-tuned on the same training set using RSLora(Kalajdzievski, 2023) with rank $r=256$ and scaling factor $\alpha=128$. Training is conducted for a single epoch with batch size 4 and learning rate $5\times 10^{-6}$, following standard practice for large-scale relevance fine-tuning. The MoE framework, including the router, projection layers, and fusion classifier, is trained with batch size 128, learning rate $1\times 10^{-4}$, and $\lambda=0.01$ in Eq. 6, while all base LLMs remain frozen.
@@ 3.4. Ablation Studies
3.5. Geometric Interpretation: Manifold Unfolding and Anisotropy
To intuitively demonstrate the effectiveness of our fusion strategy, we visualize the latent representations of the test set using t-SNE (t-Distributed Stochastic Neighbor Embedding). We randomly sampled 45,000 instances from the test set, balanced between relevant (Class 1, blue) and irrelevant (Class 0, red) samples. The visualization compares the final embedding space generated by our Concatenation-based framework (Figure  3 (a)) against the Weighted Fusion baseline (Figure  3(b)). All hyperparameters, including perplexity and learning rate, were kept consistent to ensure a fair topological comparison. This topological separation visualized in Figure  3 (a) serves as the geometric underpinning for the superior AUC scores reported in Table 2.
Figure 3. t-SNE Visualization of Embedding Spaces under Different Fusion Strategies. Comparison between (a) Concatenation and (b) Weighted Fusion. The concatenation strategy (a) yields a highly structured manifold with a clear separation gap, effectively avoiding the class entanglement observed in the weighted averaging approach (b).
The comparative visualization reveals fundamental topological differences that align with our theoretical hypothesis of Manifold Preservation, providing a geometric explanation for the performance gains reported in Table  2.
Table 3. Online performance comparison of the distilled student model (ColBERT) trained by the strongest single teacher (Sailor2) versus our Proposed MoE across six markets. ”Bad Ratio” ($\downarrow$) indicates the percentage of irrelevant results (lower is better), while ”Online AUC” ($\uparrow$) measures ranking accuracy.
Bad Ratio ($\downarrow$) AUC ($\uparrow$)
Model Overall ID MY PH SG TH VN
Sailor2-20B 9.00 10.48 8.81 9.63 7.45 8.06 9.65 90.39
Proposed MoE 8.55 10.26 8.45 9.32 6.90 7.47 8.90 91.28
Manifold Unfolding: Topological Detachment vs. Entanglement The most striking observation in Figure  3 (a) is the emergence of a distinct, isolated subspace for Class 0 samples (visible as the dense red cluster detached on the right). This phenomenon indicates that our method successfully achieves ”Manifold Unfolding.” By concatenating the representations ($\mathbf{h_{A}};\mathbf{h}_{B}]$), the system preserves the full dimensionality of the heterogeneous expert spaces. This allows the model to project conflicting samples into orthogonal subspaces, creating a ”clean gap” (white space) that makes the data linearly separable for the subsequent MLP.
Beyond topology, these results illustrate the critical role of anisotropy preservation. Pre-trained LLM embeddings are typically anisotropic, encoding semantic information in specific, high-variance directions (”spikes”). When performing weighted averaging on unaligned experts, the dominant semantic direction of one model often acts as noise to the other. This leads to destructive interference, where sharp distinctive features are canceled out, pushing the distribution towards isotropy (a more uniform, spherical distribution). The result is the visual ”smearing” seen in Figure  3(b) where the boundaries between classes are softened.
Our Concatenation strategy maintains the orthogonal anisotropy of the experts. The sharp, elongated separation seen in Figure  3(a) proves that the high-variance semantic signals from both models are preserved intact. This allows the downstream classifier to leverage the ”sharpest” features from either expert to define a precise decision boundary, effectively resolving the ”curse of multilingualism.”
@@ 3.6. Online Deployment
In industrial search advertising systems, upon receiving a user query, a Top-K subset of ads is selected from a vast inventory, following a standard multi-stage pipeline: index → retrieval → prerank → rank → auction → rerank. At the retrieval stage, the relevance scores serve dual purposes: filtering candidates for downstream stages and informing dynamic reserve prices in the auction mechanism (Dütting et al., 2017; Li, 2025), which balance user experience, platform revenue, and advertiser ROI. Accurate relevance estimation is therefore critical for both result quality and monetization efficacy.
However, industrial-scale retrieval involves tens to hundreds of thousands of candidate items per query. Real-time inference with large language models (LLMs) at this scale is computationally prohibitive due to latency constraints. To leverage LLM semantics without losing efficiency, we use offline knowledge distillation: LLMs generate high-quality relevance labels on a billion-scale query-item corpus to train a compact student model. We select ColBERT (Khattab and Zaharia, 2020) as the student model due to its strong representational capacity and efficiency in late-interaction architectures. The distilled ColBERT model is deployed online, enabling fine-grained, low-latency relevance scoring that meets the demands of real-time retrieval.
Online Evaluation. To verify whether the MoE is a better teacher than single monolithic models, we compare the performance of ColBERT students distilled from different teachers (Qwen, Gemma, Sailor, and our MoE). We focus on two key metrics:
  • Bad Ratio($\downarrow$): The percentage of irrelevant items in the top retrieval results, evaluated by human experts on a sampled traffic set. Lower is better.
  • Online AUC($\uparrow$): The ranking capability of the student model on the held-out test set.
Results. To validate the effectiveness of our framework in a production setting, we benchmark the MoE-distilled student against a student distilled from Sailor2-20B, which served as the strongest single-model baseline in our offline experiments. As presented in Table 3, the results demonstrate the comprehensive superiority of the MoE teacher:
Universal Quality Improvement: The MoE-distilled student achieves a significantly lower Overall Bad Ratio (8.55%) compared to the Sailor baseline (9.00%), representing a relative reduction of 5%. Notably, our approach outperforms Sailor across all six markets, with the most substantial gains observed in Vietnam (9.65% → 8.90%) and Thailand (8.06% → 7.47%). This confirms that the MoE’s diverse expert knowledge is effectively transferred, mitigating the performance fluctuations often seen in single models across different languages.
Enhanced Ranking Accuracy: Beyond filtering irrelevant results, the MoE teacher also imparts better ranking capabilities. The Online AUC improves from 90.39% to 91.28%, indicating that the student model has successfully learned the fine-grained semantic nuances captured by the manifold-preserving fusion.
This result validates that the MoE teacher provides cleaner, more robust supervision than single models, successfully transferring its fine-grained multilingual understanding to the deployable student model.